OpenStudy (anonymous):

A car slows down from 22 m/s to rest in a distance of 83 m. What was its acceleration, assumed constant?

OpenStudy (anonymous):

use one of the five kinematics equation \(v_f^2=v_i^2+2ad\) solving for \(a\) will give you: \(a=\frac{v_f^2-v_i^2}{2d}\) in your case \(v_f=0,\ v_i=22\, d=83\)